OpenStudy (anonymous):

Need help now for problems. Need every step to problem. _12+square root of 45 divided by 3

OpenStudy (anonymous):

is it ....(12 +sqrt(3*3*5) )/3 =3((4+sqrt(5) )/3 = 4+sqrt(5)

OpenStudy (anonymous):

so the answer choices are -4+i square root of 3, or 4+i square root of 5 or -4+i square root of 5 or -4-i square root of 5

OpenStudy (anonymous):

can u rewrite the qustions properly

OpenStudy (anonymous):

sure...........is the answer 4+sq root of 5?

OpenStudy (anonymous):

\[\frac{12+\sqrt{45}}{3}\]?

OpenStudy (anonymous):

there is no "i" in it

OpenStudy (anonymous):

\[\frac{12+\sqrt{-45}}{3}=\frac{12}{3}+\frac{3\sqrt{5}i}{3}=4+\sqrt{5}i\]